### v1

# Vibe Researching as Wolf Coming: Can AI Agents with Skills Replace or Augment Social Scientists?

Abstract
AI agents—systems that execute multi-step reasoning workflows with persistent state, tool access, and specialist skills—represent a qualitative shift from prior automation technologies in social science. Unlike chatbots that respond to isolated queries, AI agents can now read files, run code, query databases, search the web, and invoke domain-specific skills to execute entire research pipelines autonomously. This paper introduces the concept of vibe researching—the AI-era parallel to “vibe coding” (Karpathy, 2025)—and uses scholar-skill, a 21-skill plugin for Claude Code covering the full research pipeline from idea to submission, as an illustrative case. I develop a cognitive task framework that classifies research activities along two dimensions—codifiability and tacit knowledge requirement—to identify a delegation boundary that is cognitive, not sequential: it cuts through every stage of the research pipeline, not between stages. I argue that AI agents excel at speed, coverage, and methodological scaffolding but struggle with theoretical originality and tacit field knowledge. The paper concludes with an analysis of three implications for the profession—augmentation with fragile conditions, stratification risk, and a pedagogical crisis—and proposes five principles for responsible vibe researching.
Keywords: artificial intelligence, AI agents, computational social science, research automation, vibe coding, large language models
1 Introduction
When was the last time you did something in your research that only you could have done? Not ran the code. Not wrote the literature review. Not formatted the citations. Something that required your specific judgment—your field knowledge, your theoretical imagination, your sense of what matters.
The question is not rhetorical. AI systems have automated code generation since 2022 (Eloundou et al., 2024), coherent text production since 2023, and tool-using multi-step workflows since 2024 (Park et al., 2023; Wu et al., 2023). In early 2025, Andrej Karpathy coined the term “vibe coding” to describe a programming style in which the user describes what they want, the AI writes the code, and the user barely looks at it—the vibe is right, so they ship it (Karpathy, 2025). This paper names the parallel phenomenon in social science: vibe researching. The title’s “wolf coming” alludes to the Chinese fable of the boy who cried wolf—but in this version, the wolf is real. You describe your research question. The AI runs the literature review, designs the study, analyzes the data, drafts the paper, and simulates the reviewers.
The core question that structures this paper is: when AI agents can execute the research pipeline, what is the researcher’s distinctive contribution? Is the researcher still the author? The designer? Or are they becoming something more like a curator—selecting among options the AI generates?
I address this question by developing a cognitive task framework and applying it to a concrete case: scholar-skill, a 21-skill plugin for Claude Code that covers the entire social science research workflow, from initial puzzle to journal submission. The system targets publication in top journals including the American Sociological Review, the American Journal of Sociology, Demography, Science Advances, Nature Human Behaviour, and Nature Computational Science.
The argument proceeds in five steps. First, I situate the current moment within four historical waves of research automation, showing that the present wave is qualitatively different because it automates reasoning across stages, not merely execution within a single stage. Second, I describe the scholar-skill system as an illustrative case of agentic AI in social science. Third, I develop a two-dimensional framework—codifiability and tacit knowledge—that identifies a delegation boundary cutting through every stage of the research pipeline. Fourth, I assess what AI agents can and cannot do, distinguishing speed and methodological scaffolding (where AI excels) from theoretical originality and tacit field knowledge (where it struggles). Fifth, I draw out three implications: the augmentation thesis and its fragile conditions, the stratification risk of an “AI productivity premium,” and a pedagogical crisis in graduate training.
This paper contributes to a growing literature on AI and social science (Bail, 2024; Ziems et al., 2024; Zhang, 2023) by offering not an empirical test but a conceptual framework grounded in an operational system. The framework’s value lies in specifying where the human–AI boundary falls—not between pipeline stages, but within them—and in naming the conditions under which augmentation succeeds or fails.
2 Background: From Automation to Agentic AI
2.1 Four Waves of Research Automation
Social science research has been progressively automated across four waves, each targeting a different type of cognitive work (Table 1). This periodization builds on task-based frameworks in labor economics that distinguish routine from non-routine work (Autor, 2015; Acemoglu and Restrepo, 2019).
Table 1: Four Waves of Research Automation in the Social Sciences
Wave Period What was automated What remained human
1 1970s–1990s Computation (SPSS, Stata) Reasoning, interpretation
2 2000s Data collection (web, APIs) Design, interpretation
3 2010s Text analysis (NLP, ML) Theory, framing
4 2024+ Multi-step reasoning Open question
The first wave (1970s–1990s) automated computation: statistical software replaced manual calculation but left reasoning, interpretation, and design entirely human (Grimmer et al., 2022). The second wave (2000s) automated data collection: web scraping, online surveys, and administrative data linkages expanded the volume of accessible data without changing how researchers designed studies or interpreted results. The third wave (2010s) automated text analysis and pattern recognition: natural language processing, machine learning classifiers, topic models, and embedding-based methods enabled the computational analysis of text at scale (Lazer et al., 2009; Grimmer et al., 2022; Nelson, 2020; Rodriguez et al., 2023). In each wave, a rule-based execution task was handed off, but the reasoning—the interpretation, the design, the theory—remained human.
The fourth wave, beginning around 2024, is different. For the first time, the thing being automated is reasoning itself—multi-step inference, argument construction, methodological judgment. Large language models (LLMs) now generate not only text but also research designs, causal diagrams, identification strategies, and manuscript sections calibrated to specific journal norms (Bail, 2024). The question of what remains irreducibly human after Wave 4 is genuinely open.
2.2 From Chatbot to AI Agent
The distinction between a chatbot and an AI agent is architectural, not merely rhetorical. A chatbot processes a single query and returns a response. It is stateless: it has no memory across exchanges, no file access, no code execution capability, and no tool use.
An AI agent, by contrast, is multi-step, persistent, and tool-using (Park et al., 2023; Wu et al., 2023). In the context of social science research, this means the agent can: read and write files to disk; execute R, Python, or Stata code via a shell; query databases (Zotero, NHANES, IPUMS); search the web and fetch papers; spawn parallel sub-agents for independent tasks; and invoke specialist skills with domain-specific knowledge bases. This is a qualitatively different architecture that enables a qualitatively different range of tasks (Shinn et al., 2024).
The inflection from assistant to agent is the critical threshold. Prior AI tools helped researchers within a stage—write better code, classify text, check grammar. AI agents can now operate across stages, maintaining state from literature review through analysis to manuscript drafting. This cross-stage persistence is what makes “vibe researching” possible and what distinguishes the current moment from the chatbot era.
2.3 Related Work
Several recent contributions address the intersection of AI and scientific research. Wang et al. (2023) provide a comprehensive review of AI across scientific disciplines, mapping capabilities from molecular simulation to materials discovery. Lu et al. (2024) demonstrate an AI system capable of generating research ideas, writing code, running experiments, and producing full scientific papers autonomously, though with significant quality limitations. Boiko et al. (2023) show that LLM-driven agents can autonomously plan and execute chemistry experiments using robotic laboratories. In the social sciences specifically, Bail (2024) argues that generative AI can augment social science across data collection, analysis, and theory development, but warns of risks including “galaxy-brained” theorizing and reduced incentives for deep methodological training. Ziems et al. (2024) provide a systematic evaluation of LLM capabilities for computational social science tasks, finding strong performance on structured tasks but limitations on those requiring deep domain knowledge. Argyle et al. (2023) demonstrate that LLMs can simulate human survey responses with surprising fidelity, raising both methodological opportunities and validity concerns.
This paper extends this literature by offering a detailed case study of an operational system—scholar-skill—and developing a cognitive task framework that specifies the human–AI delegation boundary with greater precision than prior work. Rather than asking “can AI do social science?” in general terms, I ask: which specific cognitive tasks within each pipeline stage can be delegated, and which cannot?
3 The Scholar-Skill System: An Illustrative Case
3.1 Architecture and Scope
Scholar-skill is a plugin for Claude Code—Anthropic’s command-line interface for the Claude language model—that organizes 21 specialist AI skills into a complete research pipeline. The system is installed locally, runs on the researcher’s laptop, and accesses local files, databases, and code environments. Table 2 summarizes the seven stage-groups and their constituent skills.
Table 2: Scholar-Skill: 21 Skills Organized by Research Stage
Stage Skills
Formulation scholar-idea, scholar-lit-review, scholar-lit-review-hypothesis, scholar-hypothesis
Design scholar-design, scholar-causal
Data scholar-data, scholar-eda
Analysis scholar-analyze, scholar-compute, scholar-ling
Writing scholar-write, scholar-citation
Ethics/Safety scholar-ethics, scholar-safety
Submission scholar-journal, scholar-respond, scholar-open
Extensions scholar-grant, scholar-presentation, scholar-full-paper
Each skill is a self-contained workflow with its own knowledge base: reference documents, code templates, journal-specific norms, and quality checklists. The orchestrator—scholar-full-paper—coordinates all 21 skills across 14 phases with 18 quality gates that prevent the pipeline from advancing without meeting minimum standards.
3.2 Key Technical Capabilities
Several technical capabilities merit description because they illustrate the distance between current AI agents and prior tools.
Literature synthesis at scale. Scholar-lit-review-hypothesis queries the researcher’s Zotero library (20,000+ items) via direct SQLite access, retrieves and extracts relevant PDFs, synthesizes themes, identifies gaps, selects theoretical frameworks from 25+ candidates, and derives testable hypotheses—producing a 1,200-word literature review section calibrated to the target journal’s citation density in under three minutes.
Causal identification. Scholar-causal constructs causal DAGs with adjustment sets, selects from eight identification strategies (OLS, difference-in-differences, regression discontinuity, instrumental variables, panel fixed effects, matching, synthetic control, and causal mediation), generates working code in both R and Stata, produces assumption statements and diagnostic tests, and writes the identification argument for the Methods section.
Asset-driven writing. Scholar-write uses a three-tier knowledge graph built from 127 published papers: 32 by the author, 8 collaborator articles, and 87 top-journal exemplars. Tier 1 provides a section-snippets library organized by nine rhetorical functions. Tier 2 provides pre-extracted stylistic annotations (opening hooks, gap sentences, contribution claims, voice registers, citation density norms) for all 127 papers. Tier 3 provides full PDF access for deep reads. The system generates original prose that mimics the rhetorical structure of published papers while producing novel content—a distinction that separates structure mimicry from plagiarism.
Peer review simulation. Scholar-respond spawns three reviewer agents in parallel—a methods specialist, a theory expert, and a senior editor—each calibrated to the target journal’s priorities. The simulation produces a triage dashboard classifying concerns as MAJOR-FEASIBLE, MINOR-EASY, DISAGREE, or CONFLICT, with cross-reviewer overlaps flagged for priority attention.
Ethics and data safety. Scholar-safety pre-scans data files for personally identifiable information using local pattern matching (no data enters the AI context), while scholar-ethics generates AI use disclosures, plagiarism audits, research integrity checks, and all journal-required declarations.
3.3 What the System Is and Is Not
Scholar-skill is not an autonomous research engine. It does not independently formulate research questions, make theoretical choices, or decide which findings are interesting. Every skill produces options—candidate framings, alternative designs, draft prose—that the researcher must evaluate, select, and often revise. The system is best understood as a highly capable research assistant: one that can execute complex tasks across the pipeline but cannot replace the judgment that makes research meaningful.
This distinction is important because the same architecture that enables productive augmentation also enables uncritical delegation. A researcher who uses scholar-skill to generate five candidate research questions and selects the most theoretically interesting one is augmented. A researcher who accepts the first suggestion without evaluation is engaged in vibe researching in its most problematic form.
4 A Cognitive Task Framework for AI-Augmented Research
4.1 Two Dimensions: Codifiability and Tacit Knowledge
To move beyond generic claims about AI “replacing” or “augmenting” researchers, I propose a framework that classifies research tasks along two dimensions.
Codifiability refers to the extent to which a task can be decomposed into explicit, rule-following procedures. Literature search is highly codifiable: query databases, apply inclusion criteria, extract themes. Theory generation is weakly codifiable: it requires creative recombination of ideas in ways that resist algorithmic specification.
Tacit knowledge requirement refers to the extent to which successful task performance depends on knowledge that cannot be fully articulated—field politics, trust networks, timing intuition, editorial culture, a sense of what is “live” versus “settled” in a subfield (Collins and Evans, 2007). Polanyi’s (1966) observation that “we know more than we can tell” captures the structural limitation: AI systems operate on explicit, codified knowledge; they have no access to the tacit dimension.
4.2 Four Task Types
Crossing these dimensions yields four task types, each with a different relationship to AI automation (Figure 1).
[…]
Figure 1: Four cognitive task types in the research pipeline, classified by automation potential and tacit knowledge requirement. Type C tasks (execution) are the most automatable. Type A tasks (formulation) are the least.
Type C (Execution): High automation potential, low tacit knowledge. Running regressions, producing descriptive statistics, generating visualizations, scanning data for sensitive information. AI agents are very good here—the tasks are codifiable and the outputs are verifiable.
Type D (Communication): Medium-high automation potential, medium tacit knowledge. Drafting prose, formatting for journals, responding to reviewers, writing grant proposals. AI can draft; the researcher must judge. This is the “dangerous middle zone” where AI output looks good but subtle errors require expert eyes (Dell’Acqua et al., 2023).
Type B (Planning): Medium automation potential, high tacit knowledge. Selecting identification strategies, constructing causal diagrams, designing open science protocols, making ethical judgments. The AI generates options; the researcher must own the choices.
Type A (Formulation): Low automation potential, very high tacit knowledge. Identifying new research questions, recognizing when existing frameworks are inadequate, importing framings from non-adjacent fields. The AI assists; the judgment is irreducibly human.
4.3 The Delegation Boundary Is Cognitive, Not Sequential
The framework reveals a critical insight: the boundary between what to delegate and what to protect is cognitive, not sequential. It does not fall between pipeline stages (“design is human, analysis is AI”). It cuts through every stage. At every stage, some tasks are codifiable and delegable; others require tacit judgment and are not. Table 3 illustrates.
Table 3: The Delegation Framework: What to Delegate, What to Protect
Task Codifiable? Tacit? Delegate?
Literature synthesis High Low Yes
Method selection & code High Medium Yes, with oversight
Analysis execution High Low Yes
Prose drafting Medium Medium Partially
RQ formulation Low High AI assists only
Theory generation Low Very high No
Field judgment None Extreme Never
The researcher who understands this framework is augmented—delegating execution while protecting judgment. The researcher who delegates everything is producing, at best, competent research with someone else’s judgment. At worst, they are publishing errors they cannot detect.
5 What AI Agents Can and Cannot Do
5.1 AI Excels at Speed and Coverage
The speed advantage is substantial. A human researcher typically spends two to three weeks reading and synthesizing 100–200 papers on a topic, shaped by what they recall and biased toward familiar journals. Scholar-lit-review-hypothesis queries a 20,000-item Zotero library in seconds, produces thematic synthesis in under ten minutes, calibrates citation density to journal norms, and does so without fatigue or availability bias.
The productivity effects of generative AI on knowledge work are empirically documented. Noy and Zhang (2023) find that ChatGPT access substantially increased writing productivity in a randomized experiment, with the largest gains for lower-ability workers. Liang et al. (2024) estimate that by early 2024, at least 10–17% of computer science papers showed evidence of LLM-assisted writing. The trajectory is clear: AI-assisted research production is already widespread and growing.
The important caveat: coverage is not comprehension. The AI synthesizes what is in the database. It does not know what matters. It does not know which findings are credible, and it introduces systematic risks when used for tasks like annotation that require cultural and contextual judgment (Lin and Zhang, 2025). It cannot sense the collective epistemic judgment of a field—what Evans and Foster (2011) call “metaknowledge,” the understanding a scientific community holds beyond the sum of individual papers.
5.2 AI Excels at Methodological Scaffolding
A researcher trained in OLS and logistic regression can, with AI scaffolding, access working code for staggered difference-in-differences with Callaway and Sant’Anna (2021) estimators, exponential random graph models, causal forests with SHAP explanations, structural topic models, and zero-shot image classification with DINOv2 or CLIP. The AI provides working code, assumption statements, diagnostics, and write-up templates.
This scaffolding is genuinely valuable. It expands the set of methods a researcher can attempt. But attempting a method and understanding it are different things. The researcher must still understand the assumptions well enough to defend them in peer review. A staggered DiD specification generated by scholar-causal is only as good as the researcher’s judgment about whether the parallel trends assumption holds in their specific empirical context. That judgment is not in the code.
5.3 AI Struggles with Theoretical Originality
AI-generated theory looks like this: “Drawing on Bourdieu’s concept of field, I argue that X shapes Y through Z mechanism, moderated by Q.” It is well-structured, correctly cited, and the mechanism is plausible. The problem: it is derived from existing frameworks. It recombines what is already there.
What AI cannot do: identify a genuinely new mechanism; recognize when existing frameworks are inadequate; import a framing from a non-adjacent field in a surprising way; make the conceptual leap that changes a subfield. Consider Granovetter (1973)’s “strength of weak ties”—the argument was not derivable from the prior network literature. Tilly (1998)’s categorical pair mechanism for durable inequality was not sitting in the literature waiting to be synthesized. These examples illustrate the irreducibly human zone that AI cannot currently enter.
Si et al. (2024) find in a large-scale evaluation that LLM-generated research ideas were rated as more novel than human-generated ones on average but scored lower on feasibility (see also Girotra et al., 2023). This finding complicates a simple narrative—AI may generate surprising combinations—but the novelty measured in that study is recombinative, not the deep theoretical innovation that characterizes paradigm-shifting work. The AI is an excellent theory articulator; it is not a theory creator. For most of what we publish, articulation is enough. For the papers that matter most, creation is what counts.
5.4 AI Struggles with Tacit Field Knowledge
The structural divide is between explicit knowledge (published literature, methods recipes, journal norms, data sources) and tacit knowledge (field politics, trust networks, timing intuition, editorial culture). The AI has the former in massive quantity. It has none of the latter.
Concretely: suppose you are writing about residential segregation. The AI can produce a competent research agenda, grounded in the literature, with sensible hypotheses. What it cannot tell you: whether the leading scholars in that subfield find your specific angle interesting or tired; whether this is a moment when segregation measurement is contested in ways that make your design politically fraught; whether a particular editor is sympathetic to structural versus behavioral arguments. That knowledge is embodied, relational, and accumulated through years of conference conversations and rejected papers (Collins and Evans, 2007).
6 Implications for the Social Sciences
6.1 The Augmentation Thesis and Its Fragile Condition
In practice, AI augmentation looks like this: a researcher who previously mastered one or two methods can now, with scaffolding, access five or more; multiple datasets can be auto-fetched and explored simultaneously; the pipeline from idea to submission operates two to three times faster; reviewer responses take days instead of months.
Augmentation works when the researcher remains the author, the judge, the intellectual owner. The AI is the research assistant, not the researcher. The human maintains the capacity for critical oversight. But this condition is fragile. Dell’Acqua et al. (2023) demonstrate the danger in a field experiment with management consultants: GPT-4 increased performance on tasks inside the AI’s capability frontier but decreased performance on tasks outside it, because consultants over-relied on AI for tasks it could not do well. The “jagged technological frontier” they describe—where AI capability boundaries are unpredictable and non-obvious—applies directly to social science research.
The question I would leave with the reader: if you never actually do the task—never run the DiD yourself, never draft the literature review by hand, never code the interviews—can you still evaluate whether the AI did it correctly? I believe the answer is no. And this is not a hypothetical future concern; it is a real pedagogical challenge right now.
6.2 The Stratification Risk
The productivity gains from AI tools are real. They are also unequally distributed. Four axes of stratification deserve attention.
Cost. Claude Pro is twenty dollars a month. API tokens for a full 14-phase pipeline run cost five to fifteen dollars. For researchers at resource-poor institutions, these are meaningful barriers.
Language. Training data skews English. Journal calibration targets English-language top journals. Non-English scholarship is disadvantaged by design.
Technical skill. Command-line setup, prompt engineering, and R/Python integration represent real barriers for researchers without computational training.
Field. The tool is calibrated to top-five sociology, political science, and computational social science journals. Area studies scholars, qualitative researchers, and Global South researchers whose work targets different venues get less value and sometimes actively incorrect guidance.
The “AI productivity premium”—the widening gap between researchers with AI access and skills and those without—is itself a stratification phenomenon, one that social scientists are well positioned to study (Acemoglu, 2024; Eloundou et al., 2024; Webb, 2020). It is also a call for design choices that prioritize access: open-weight models, documented prompts, and shared tools.
6.3 The Pedagogical Crisis
Traditional PhD training teaches students to run regressions, code interviews, write literature reviews, and respond to reviewers. These are execution skills. AI agents can now execute all of them. This does not make these skills worthless—it changes what they are for.
A student who understands difference-in-differences at a deep level—who has read Callaway and Sant’Anna, who understands the parallel trends assumption well enough to test it, who has manually worked through a staggered adoption case—can evaluate AI-generated DiD code. They can see when the parallel trends test is being misapplied. A student who only knows “run the DiD package” cannot.
The risk: if we train students to operate pipelines, and AI can operate pipelines, we are training them for a depreciating skill. The response is not to ban AI tools. It is to teach methods as foundations for evaluation, not just production. And it is to invest more heavily, not less, in deep theory training—because theory is precisely where AI is weakest and human scholars are irreplaceable (Bail, 2024).
7 Toward Responsible Vibe Researching
I propose five principles for responsible use of AI agents in social science research.
Disclose. Report AI assistance in the methods section: what was used, when it was used, and which parts of the paper it contributed to. Example: “Sections 3–5 were initially drafted with AI assistance using scholar-write; all content was reviewed and revised by the author.” This is not a confession. It is a practice we need to normalize.
Verify. Review all AI-generated code, analysis, and prose before publishing. Run the code yourself. Check the citations. Errors published under your name are yours. Scholar-skill includes verification sub-agents that audit code and results, but these supplement—never replace—human oversight.
Maintain skills. Deliberately practice the tasks you delegate. Run regressions by hand sometimes. Write a section without AI sometimes. Not because AI assistance is cheating, but because oversight capacity depends on practiced judgment. The researcher who never does the work can no longer evaluate whether the AI did it correctly.
Protect originality. The research question and theoretical contribution must remain the researcher’s own. That is not just a norm—it is where the intellectual value of the work lives. AI generates options; the researcher’s theoretical imagination and field knowledge determine which options are worth pursuing.
Design for access. Use open models where feasible. Document prompts. Share tools. Do not let the AI productivity premium become another axis of academic stratification. The differential adoption of AI tools across institutions, languages, and fields is a social inequality that researchers can both study and mitigate through design choices.
8 Discussion and Conclusion
This paper has made three arguments.
First, AI agents are qualitatively different from prior tools. They execute multi-step reasoning workflows—not just computation or text—across the entire research pipeline. The scholar-skill system demonstrates that a single plugin can now cover 21 distinct research tasks from idea formalization to journal submission, coordinated by an orchestrator with quality gates. The change in what a solo researcher can accomplish is real and significant.
Second, the delegation boundary is cognitive, not sequential. It cuts through every stage. At every point in the pipeline, some tasks are codifiable and delegable while others require tacit judgment and are not. The framework developed here—classifying tasks by codifiability and tacit knowledge requirement—provides a practical tool for researchers deciding what to delegate. The rule is simple: delegate codifiable execution; protect tacit judgment.
Third, vibe researching is already here; the profession’s normative response is not. Liang et al. (2024) document the rapid spread of LLM-assisted writing in scientific papers. The systems are operational and improving. Disclosure norms, pedagogical reform, equitable design, and deeper theory training are the four urgent interventions. We are not waiting for a future that may arrive. We are in it.
Several limitations warrant acknowledgment. The case study is based on a single system (scholar-skill) developed by the author, which may not generalize to all AI research tools. The cognitive task framework, while grounded in the operational characteristics of the system, has not been empirically validated through user studies or controlled experiments. The four-type taxonomy simplifies a continuous space; some tasks may resist clean classification. Future work should empirically test the framework’s predictions about delegation effectiveness, conduct user studies comparing augmented and unaugmented research workflows, and examine how AI tool adoption varies across disciplines, institutions, and career stages.
The wolf is at the door. The question is not whether to let it in—it is already in. The question is what role social scientists play, and what they protect. AI may transform social science research before social scientists have studied how AI transforms social science research. Our task is to use AI to amplify our reach while protecting our capacity for original inquiry—and to study AI itself as a social phenomenon shaping labor, inequality, and knowledge production.

### v2

@@
-Vibe Researching as Wolf Coming: Can AI Agents with Skills Replace or Augment Social Scientists?
+arXiv:2602.22401v2[cs.AI] 27 Feb 2026
 Abstract
@@
 Unlike chatbots that respond to isolated queries, AI agents can now read files, run code, query databases, search the web, and invoke domain-specific skills to execute entire research pipelines autonomously.
-This paper introduces the concept of vibe researching—the AI-era parallel to “vibe coding” (Karpathy, 2025)—and uses scholar-skill, a 21-skill plugin for Claude Code covering the full research pipeline from idea to submission, as an illustrative case.
+This paper introduces the concept of vibe researching—the AI-era parallel to “vibe coding” (Karpathy, 2025)—and uses scholar-skill, a 23-skill plugin for Claude Code covering the full research pipeline from idea to submission, as an illustrative case.
 I develop a cognitive task framework that classifies research activities along two dimensions—codifiability and tacit knowledge requirement—to identify a delegation boundary that is cognitive, not sequential: it cuts through every stage of the research pipeline, not between stages.
@@
 Or are they becoming something more like a curator—selecting among options the AI generates?
-I address this question by developing a cognitive task framework and applying it to a concrete case: scholar-skill, a 21-skill plugin for Claude Code that covers the entire social science research workflow, from initial puzzle to journal submission.
+I address this question by developing a cognitive task framework and applying it to a concrete case: scholar-skill, a 23-skill plugin for Claude Code that covers the entire social science research workflow, from initial puzzle to journal submission.
 The system targets publication in top journals including the American Sociological Review, the American Journal of Sociology, Demography, Science Advances, Nature Human Behaviour, and Nature Computational Science.
@@
 In the context of social science research, this means the agent can: read and write files to disk; execute R, Python, or Stata code via a shell; query databases (Zotero, NHANES, IPUMS); search the web and fetch papers; spawn parallel sub-agents for independent tasks; and invoke specialist skills with domain-specific knowledge bases.
-This is a qualitatively different architecture that enables a qualitatively different range of tasks (Shinn et al., 2024).
+This is a qualitatively different architecture that enables a qualitatively different range of tasks (Shinn et al., 2023).
 The inflection from assistant to agent is the critical threshold.
@@
 3.1 Architecture and Scope
-Scholar-skill is a plugin for Claude Code—Anthropic’s command-line interface for the Claude language model—that organizes 21 specialist AI skills into a complete research pipeline.
+Scholar-skill is a plugin for Claude Code—Anthropic’s command-line interface for the Claude language model—that organizes 23 specialist AI skills into a complete research pipeline.
 The system is installed locally, runs on the researcher’s laptop, and accesses local files, databases, and code environments.
-Table 2 summarizes the seven stage-groups and their constituent skills.
-Table 2: Scholar-Skill: 21 Skills Organized by Research Stage
+Table 2 summarizes the ten stage-groups and their constituent skills.
+Table 2: Scholar-Skill: 23 Skills Organized by Research Stage
 Stage Skills
@@
 Submission scholar-journal, scholar-respond, scholar-open
+Replication scholar-replication
+Quality Assurance scholar-auto-improve
 Extensions scholar-grant, scholar-presentation, scholar-full-paper
 Each skill is a self-contained workflow with its own knowledge base: reference documents, code templates, journal-specific norms, and quality checklists.
-The orchestrator—scholar-full-paper—coordinates all 21 skills across 14 phases with 18 quality gates that prevent the pipeline from advancing without meeting minimum standards.
+The orchestrator—scholar-full-paper—coordinates all 23 skills across 17 phases with 43 quality gate items and three hard stops (at data safety, citation verification, and ethics compliance) that prevent the pipeline from advancing without meeting minimum standards.
 3.2 Key Technical Capabilities
 Several technical capabilities merit description because they illustrate the distance between current AI agents and prior tools.
+Idea formalization.
+Scholar-idea converts a broad research intuition into a formal, testable research question through a 10-step workflow: puzzle reframing, multi-angle exploration, variable mapping, hypothesis derivation, feasibility and novelty assessment, and a five-agent multi-agent evaluation panel (methodologist, theorist, domain expert, skeptic, and integrator) that stress-tests the proposed research question before the researcher commits to a design.
 Literature synthesis at scale.
@@
 The system generates original prose that mimics the rhetorical structure of published papers while producing novel content—a distinction that separates structure mimicry from plagiarism.
+Each section undergoes a five-agent internal review panel—evaluating logic, rhetoric, journal fit, citation evidence, and accessibility—followed by a synthesizer that consolidates critiques with cross-agent agreement scoring and a reviser that implements changes, with user confirmation before advancing.
 Peer review simulation.
-Scholar-respond spawns three reviewer agents in parallel—a methods specialist, a theory expert, and a senior editor—each calibrated to the target journal’s priorities.
+Scholar-respond spawns three or four reviewer agents in parallel—a methods specialist, a theory expert, a senior editor, and for computational papers, a fourth computational methods reviewer—each calibrated to the target journal’s priorities.
 The simulation produces a triage dashboard classifying concerns as MAJOR-FEASIBLE, MINOR-EASY, DISAGREE, or CONFLICT, with cross-reviewer overlaps flagged for priority attention.
@@
 Scholar-safety pre-scans data files for personally identifiable information using local pattern matching (no data enters the AI context), while scholar-ethics generates AI use disclosures, plagiarism audits, research integrity checks, and all journal-required declarations.
+Replication package construction.
+Scholar-replication builds, documents, tests, and verifies a journal-ready replication package.
+It assembles scripts from the analysis pipeline into a structured directory with a master execution script, generates an AEA-template README with nine sections, runs clean-run validation in an isolated environment comparing reproduced outputs against originals, performs a paper-to-code correspondence audit mapping every table, figure, and in-text statistic to its producing script, and prepares archive deposits for Zenodo or Harvard Dataverse with DOI assignment.
+This distinguishes it from scholar-open, which produces open science declarations (templates, checklists); scholar-replication actively builds the package.
+Continuous quality improvement.
+Scholar-auto-improve provides a self-improving quality engine operating in four modes: observe (monitor skill executions and log performance patterns), audit (systematically evaluate skill outputs against quality checklists), improve (propose and implement refinements to skill instructions, reference files, and quality gates), and evolve (identify gaps in the skill ecosystem and generate new capabilities).
+This meta-skill enables the system to learn from each research engagement, progressively tightening quality gates and expanding coverage—a capacity absent from static tool pipelines.
+Multi-format manuscript output.
+The system produces manuscripts in four formats—Markdown, Word (.docx), , and PDF—via pandoc conversion, ensuring compatibility with diverse journal submission systems and co-author workflows.
 3.3 What the System Is and Is Not
@@
 The productivity gains from AI tools are real. They are also unequally distributed. Four axes of stratification deserve attention.
-Cost. Claude Pro is twenty dollars a month. API tokens for a full 14-phase pipeline run cost five to fifteen dollars. For researchers at resource-poor institutions, these are meaningful barriers.
+Cost. Claude Pro is twenty dollars a month. API tokens for a full 17-phase pipeline run cost five to fifteen dollars. For researchers at resource-poor institutions, these are meaningful barriers.
 Language. Training data skews English. Journal calibration targets English-language top journals. Non-English scholarship is disadvantaged by design.
@@
 They execute multi-step reasoning workflows—not just computation or text—across the entire research pipeline.
-The scholar-skill system demonstrates that a single plugin can now cover 21 distinct research tasks from idea formalization to journal submission, coordinated by an orchestrator with quality gates.
+The scholar-skill system demonstrates that a single plugin can now cover 23 distinct research tasks from idea formalization to journal submission, coordinated by an orchestrator with quality gates and hard stops.
 The change in what a solo researcher can accomplish is real and significant.